a panda can eat the eucalyptus leaves from a branch in 14 minutes. together, two pandas can eat the leaves from that same branch in 9 minutes. how long would it have taken the second panda to eat the leaves from the branch by itself?
Found 2 solutions by checkley77, richwmiller:
Answer by checkley77(12844)

There are 2 different ways to set up this type of problem.
Let x & y be the 2 independent times & z be the combined time working together.
The product of the independent times (xy) divided by the sum of the independent times (x+y) equals the combined time (z).
x*y/(x+y)=z
OR
1 divided by one of the independent times (1/x) plus 1 divided by the second independent times (1/y) equals 1 divided by the combined time (1/z).
1/x+1/y=1/z
both formulas are the same -- they are just presented in different forms.
1/14+1/x=1/9
1/x=1/9-1/14
1/x=(14-9)/9*14
1/x=5/126
5x=126
x=126/5
x=25.2 minutes.

Answer by richwmiller(17219)

Koalas (from Australia) eat eucalyptus (from
Australia). Pandas from China eat bamboo (from
China).